OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (texaschic101):

y - 4 = 3/4(x + 8) --- distribute through the parenthesis y - 4 = 3/4x + 6 --- add 4 to both sides y = 3/4x + 6 + 4 y = 3/4x + 10 --- multiply by the LCD, which is 4 4y = 3x + 40 --- subtract 3x from both sides -3x + 4y = 40 --- multiply by -1 to make x positive 3x - 4y = -40